OpenStudy (anonymous):

Two lines, A and B, are represented by the equations given below: Line A: 2x + 2y = 8 Line B: x + y = 4 Which statement is true about the solution to the set of equations? It is (8, 4). It is (4, 8). There is no solution. There are infinitely many solutions.

OpenStudy (kidrah69):

Well you can use elmination to solve this. \[-2( x + y = 4)\] you want opposite signs so that the numbers will cancel out

OpenStudy (kidrah69):

solve that :)

OpenStudy (anonymous):

I don't know how

OpenStudy (kidrah69):

just distribute the -2 to the whole equation

OpenStudy (anonymous):

-2 as x and y ?

OpenStudy (kidrah69):

yep and now -2 x 4=?

OpenStudy (anonymous):

-8

OpenStudy (kidrah69):

\[\frac{ 2x + 2y = 8 }{ -2x-2y=-8}\] when you subtract all of this what do you get? this shouldnt be too hard :)

OpenStudy (kidrah69):

0=0 infinte

OpenStudy (anonymous):

x -y = -8

OpenStudy (anonymous):

is that right ?

OpenStudy (kidrah69):

nope

OpenStudy (kidrah69):

the 2s cancel out so 0 and -8-8=0 therefore you get 0=0 that means infinte solutions